Write an equation in standard form for a line passing through the pair of points.
(2, -2) and (0, 3)

Answer by nerdybill(7384)

Write an equation in standard form for a line passing through the pair of points.
(2, -2) and (0, 3)
.
First, find the slope between the two points:
m (slope) = (y2-y1)/(x2-x1)
m (slope) = (3-(-2))/(0-2)
m (slope) = (3+2)/(0-2)
m (slope) = (5)/(-2)
m (slope) = -5/2
.
Using the slope above and one point (0,3) plug it into the point slope form:
y-y1 = m(x-x1)
y-3 = -5/2(x-0)
y-3 = -5/2(x)
5/2(x) + y - 3 = 0 (this is what they're looking for)